# ReDeLEx: A Framework for Relational Deep Learning Exploration

### v1

@@ 4.0.1 Databases
To establish a comprehensive yet manageable list from the overall 70+ available RDBs [29] for the RDL experimentation, we separated databases that exhibit the tabular (Sec. 3.2.2) or graph-like (Sec. 3.2.3) characteristics, or are artificially555with a single exception of the tpcd database created (see App. Table 6 for the most suitable databases).
@@ 4.0.2 RDL Models
    GraphSAGE with Linear Transformation is the simplest of the RDL models, applying a linear transformation on top of a concatenation of the attribute $a_{1},\dots,a_{n}$ embeddings $h_{v}^{(0)}\in\mathbb{R}^{n\cdot d^{(0)}}$ to yield a single embedding vector $h_{v}^{(1)}={W}h_{v}^{(0)}$ for each node $v$. The projected node embeddings $h^{(1)}\in\mathbb{R}^{d_{\phi(v)}}$ then form input into the GraphSAGE [18] model, forming the GNN stage. Finally, a task specific model head is applied.
@@
    GraphSAGE with Tabular ResNet is similar to the previous, with the tabular-level stage reducing the node embedding dimensionality, however, the operation is performed through a more sophisticated tabular ResNet model [17]. Notably, this model was previously used in [33], allowing to directly align results between the RelBench and ReDeLEx benchmarks.
@@
    DBFormer is an implementation of the Transformer-based RDL model from [32]. In constrast to the previous, [^6] The key difference can be viewed analogously to the “fusion” and “cooperation” in multi-modal learning [20, 27], considering the attributes as modalities. While the first two models fuse the representations of the discrete attributes at the beginning, the DBFormer allows for cooperation of the attributes through the GNN stage. the model retains the original node embedding dimensionality $h_{v}^{(l)}\in\mathbb{R}^{n\times d^{(l)}}$ while exploiting the attention mechanism [37] for learning interactions between both the attributes and tuples through a custom message-passing scheme.
@@ 4.0.3 Classical Models
In addition to the selected RDL models, we include key representatives from related ML domains, including Gradient Boosted Decision Trees (GBDT; [31]), Deep Tabular Learning (DTL; [3]) and Propositionalization (Prop.; [24]). Particularly, we compare against the LightGBM [22]—representative of GBDT; the getML’s [11] FastProp feature generator combined with XGBoost [8]—representative of propositionalization; and the standalone tabular ResNet [17]—representative of deep tabular learning. Importantly, the LightGBM and the ResNet have access only to data from the task’s target table, as these models fall into the tabular learning category. In contrast, the propositionalization method of FastProp with XGBoost exploits the full RDB structure.
@@
Binary Classification
ergastf1 orig. .579 .605 .086 .081 .543 .590 .910 .904 .913 .907 .914 .910
expend. orig. .852 .856 .295 .294 .846 .847 .920 .918 .918 .915 .894 .892
geneea orig. .994 .993 .317 .258 .990 .983 .985 .976 .986 .981 .996 .995
temp. .997 .990 .438 .455 .989 .982 .984 .971 .989 .973 .998 .991
hepatitis orig. .666 .626 .044 .055 .716 .634 1.0 .997 1.0 1.0 1.0 .996
imdb orig. .986 .986 .451 .454 .985 .985 .993 .993 .993 .993 .993 .993
mondial orig. .500 .500 NaN NaN .500 .542 .987 .988 1.0 .921 .987 .932
movielens orig. .583 .611 .223 .211 .563 .609 .790 .798 .801 .809 .778 .813
musk L orig. .500 .500 .381 .240 .643 .700 .905 .587 .937 .547 .984 .720
musk S orig. .500 .500 .200 .222 .625 .583 .867 .889 .933 .889 .800 .759
mutagen. orig. .917 .821 .095 .024 .917 .940 .952 .806 .964 .798 .960 .762
ncaa orig. .687 .426 .335 .212 .692 .559 .848 .759 .897 .801 .826 .720
stud.loan orig. .500 .500 .272 .247 .506 .532 1.0 1.0 1.0 1.0 1.0 1.0
amazon ichurn .738 .740 NaN NaN .653 .654 .823 .826 .823 .827 .818 .821
uchurn .579 .579 NaN NaN .526 .531 .707 .707 .707 .707 .703 .704
avito clicks .565 .539 NaN NaN .540 .541 .653 .635 .654 .672 .654 .649
visits .536 .528 NaN NaN .520 .516 .702 .662 .702 .665 .699 .660
f1 dnf .684 .685 .324 .230 .683 .708 .723 .757 .744 .740 .733 .759
top3 .729 .728 .300 .158 .661 .733 .792 .844 .714 .832 .766 .835
stack badge .785 .761 .083 .097 .689 .676 .901 .890 .901 .891 .897 .885
engm. .842 .832 .070 .071 .788 .782 .902 .906 .903 .907 .902 .903
trial sout .676 .711 .032 .034 .624 .664 .662 .702 .655 .692 .645 .675
Multiclass Classification
accidents orig. .379 .370 .506 .505 .363 .359 1.0 1.0 1.0 1.0 1.0 1.0
temp. .186 .170 .325 .336 .204 .187 .819 .583 .805 .566 .825 .727
craftbeer orig. .411 .266 .003 .017 .505 .231 .381 .255 .488 .329 .513 .356
dallas orig. .594 .581 NaN NaN .585 .572 .547 .499 .561 .473 .688 .511
temp. .570 .584 .288 .512 .680 .247 .589 .393 .615 .424 .580 .513
diabetes orig. .190 .190 .402 .383 .190 .190 .878 .870 .892 .883 .890 .875
financial orig. .461 .458 .899 .658 .698 .470 .586 .490 .738 .729 .644 .539
temp. .495 .492 .636 .632 .456 .449 .495 .492 .538 .340 .553 .256
genes orig. .060 .060 .066 .100 .060 .060 1.0 .898 1.0 1.0 1.0 .927
hockey orig. .707 .721 .395 .404 .707 .703 .714 .728 .714 .724 .713 .740
legalacts orig. .933 .924 .320 .316 .837 .824 .834 .822 .837 .825 .831 .821
temp. .870 .851 .226 .220 .788 .688 .791 .721 .786 .698 .784 .703
premiere- orig. .380 .411 .641 .601 .589 .362 .659 .400 .593 .332 .571 .352
league temp. .317 .222 .647 .639 .615 .370 .650 .384 .619 .390 .556 .407
tpcd orig. .191 .182 NaN NaN .191 .185 .733 .732 .546 .545 .722 .728
webkp orig. .139 .130 .350 .366 .139 .130 .315 .193 .295 .212 .355 .282
Avg. Rank 3.88 4.91 4.38 2.67 2.51 2.47
2.43 3.28 2.8 1.36
@@ 4.1 Benchmarking tasks
We present comprehensive results over two types of node-level tasks—binary classification and multiclass classification. The tasks can be further differentiated by the origin of the target labels and usage of temporal values. Tasks performed on datasets from the RelBench collection use generated target table attributes, while tasks on datasets from the CTU Relational use existing target table attributes. Additionally, tasks from the CTU Relational can be both static and temporal, while tasks from the RelBench collection are always temporal. Static and temporal tasks differentiate based on the constrains forced upon the sampling algorithm while generating a sub-subgraph used for training the model, and by the method of splitting the dataset between the training, validation and test data. Static tasks use neighborhood sampling [18] constrained only by the maximum number of neighbors, and data splitting is carried out at random w.r.t. a given ratio (e.g. 70:15:15). In contrast, temporal tasks extend the neighborhood sampling by incorporating temporal constraints, assuming only directed edges from nodes with an older timestamp and, similarly, the splits are carried out w.r.t. the timestamps where all training entities must precede validation and testing data, respectively (see App. 0.A for the full experimental setup).
The results on classification tasks shown in Table 1 demonstrate a strong performance of the RDL models (Sec. 4.0.2) over the classical models (Sec. 4.0.3) on majority of the datasets. Specifically, the ResNet SAGE performs very well on binary classification tasks, especially on tasks with a training table generated from historical data. The higher complexity of the DBFormer somewhat surprisingly seems to lead to higher robustness as the model slightly outperforms the ResNet SAGE in the average rank. Nevertheless, in a couple of cases, the simplest Linear SAGE model matches or even outperforms the score of both the models. Notably, on the hepatits, mondial, student loan, accidents, and genes datasets, all the RDL models present near-perfect predictions while the classical models show an order of magnitude worse score, highlighting the contribution of the RDL representation.
@@
Base Joined Base Joined 2 layers Best
@@
ergastf1 orig. .579 .605 .919 .904 .543 .590 .899 .892 .909 .905 .910 .904
hepatitis orig. .666 .626 .940 .913 .716 .634 .718 .632 1.0 1.0 1.0 1.0
mondial orig. .500 .500 1.0 .815 .500 .542 1.0 .898 .988 .944 .988 .944
movielens orig. .583 .611 .574 .622 .563 .609 .563 .609 .688 .691 .784 .796
ncaa orig. .687 .426 .797 .776 .692 .559 .819 .741 .801 .728 .861 .764
studentloan orig. .500 .500 1.0 1.0 .506 .532 1.0 1.0 1.0 1.0 1.0 1.0
amazon uchurn .579 .579 .577 .577 .526 .531 .526 .531 .705 .705 .706 .707
avito clicks .565 .539 .564 .556 .540 .541 .540 .541 .647 .658 .653 .649
visits .536 .528 .536 .532 .520 .516 .520 .516 .690 .656 .702 .664
stack badge .785 .761 .783 .759 .689 .676 .689 .676 .898 .887 .901 .891
@@
accidents orig. .379 .370 .377 .372 .363 .359 .366 .365 1.0 1.0 1.0 1.0
temp. .186 .170 .177 .163 .204 .187 .198 .178 .790 .579 .798 .582
craftbeer orig. .411 .266 .488 .245 .505 .231 .505 .231 .442 .301 .444 .311
diabetes orig. .190 .190 .190 .190 .190 .190 .190 .190 .885 .876 .885 .876
genes orig. .060 .060 .060 .060 .060 .060 .060 .060 .989 .922 1.0 .896
premiere- orig. .380 .411 .470 .416 .589 .362 .584 .325 .603 .371 .603 .371
league temp. .317 .222 .527 .485 .615 .370 .598 .399 .616 .431 .616 .431
tpcd orig. .191 .182 .846 .841 .191 .185 .676 .666 .579 .582 .579 .582
webkp orig. .139 .130 .139 .130 .139 .130 .139 .130 .265 .185 .308 .221
@@ 4.2 Tabular Learning
In this scenario, we compare results of the Tabular Learning (TL) models from the previous section to ones trained on new tables formed by join operations over the RDBs (Sec. 3.2.2). Particularly, we evaluate the TL models on tables generated by joins over both one-to-one and many-to-one relationships of the target table. [^7] This is similar to a recent RelGNN method [7], albeit limited to the target table. Additionally, we include evaluation of RDL models with exactly $2$ layers in the graph neural stage, which is conceptually equivalent to the join operation. Finally, we include the overall best RDL models to put the results into context. Note that the previous Table 1 demonstrated that the TL methods perform significantly worse on an absolute majority of tasks. In this experiment, we aim to assess whether a simple RDB transformation could actually change the situation in some cases. Particularly, we select a subset of tasks from Table 1 where both the TL models showed at least $0.1$ worse score than the best RDL model. The results in Table 2 show that, indeed, on a number of datasets the TL models register a significant improvement with results sometimes comparable to the best of RDL. Notably, on the ncaa, premierleague and tpcd they even set new best results. This experiment demonstrates existing weak spots in the new RDL approach [15], suggesting that caution and thorough analysis are still in order before deploying RDL on an RDB task.
@@ 4.3 Essential Characteristics
#Tab. 6 3 7 8 3 4 8 3 4 13.5
#FK 6 2.25 6 12 2.5 5 8 3 4 14
#Factual 35 14 24 45 17 47 110 27 40 119
#Cat. 11 2.25 11 16 6 17 44 9 14 59
#Num. 6 2 4.5 13 4 6 12 2.5 3 9
#Text 6 2 6 10 1 2 3.5 3 14 20
#Time 2 0 2 7 0.5 3 5 1.5 3 5
#Rows 472k 96k 1.36M 5.59M 16.6k 81.8k 1.6M 32.1k 804k 2.54M
#Links 752k 182k 2.22M 7.98M 32.5k 82.6k 1.58M 45.6 851k 3.49M
Diameter 3 2 2.5 3 2 2 3 2 3 4.5
1-to-1 0 0 0 2 0 0 4 0 2 3
1-to-M 5 2 4.5 11 2.5 5 5 2.5 4 12.5
#Train 12k 2.22k 27.1k 1.15M 305 546 21.8k 1.96k 18.5k 348k
#T.Factual 5 1 3 6 2.5 5 5 4 10 15
#T.Cat 2 0 0 2 0.5 1 4 2 4 10
#T.Num 0 0 0 1 0 0 2 0 1 2
#T.Text 1 0 1 3 0 0 1.5 0 1 5.5
#T.Time 1 0 0 1 0 1 1 1 1 2
Eccentric. 10.6 7.13 15.3 19.5 3.3 9.7 15.8 3.39 10.6 17.4
Density 0.025 0.001 0.018 0.079 0.004 0.028 0.102 0.003 0.059 0.118
@@
The overarching aim of ReDeLEx is to assess common characteristics of RDBs and tasks in the context of a used learning approach. While a fully comprehensive assessment is out of scope of this short paper, in Table 3 we summarize characteristics of the databases against the respective performances of the various model types. Following the analysis, RDL models generally tend to perform well on datasets with a large number of training samples and links. Propositionalization achieves best results mostly on smaller datasets with a low number of one-to-many relationships, yet with a large number of factual (non-key) columns. This is in line with some previous studies [26], despite the remaining prevalence of propositionalization methods in practice [11]. The TL methods then tend to perform best when there is a higher number of factual columns in the target table, which aligns with natural intuition. Moreover, these allow capturing more diverse attribute types where, e.g., both LightGBM and deep TL models are capable of utilizing textual attributes.

### v2

@@ 4.0.1 Databases
To establish a comprehensive yet manageable list from the overall 70+ available RDBs [motl2015ctu] for the RDL experimentation, we separated databases that exhibit the tabular (Sec. 3.2.2) or graph-like (Sec. 3.2.3) characteristics, or are artificially555with a single exception of the tpcd database created (see App. Table 6 for the most suitable databases).
@@ 4.0.2 RDL Models
    GraphSAGE with Linear Transformation is the simplest of the RDL models, applying a linear transformation on top of a concatenation of the attribute $a_{1},\dots,a_{n}$ embeddings $h_{v}^{(0)}\in\mathbb{R}^{n\cdot d^{(0)}}$ to yield a single embedding vector $h_{v}^{(1)}={W}h_{v}^{(0)}$ for each node $v$. The projected node embeddings $h^{(1)}\in\mathbb{R}^{d_{\phi(v)}}$ then form input into the GraphSAGE [hamilton2017inductive] model, forming the GNN stage. Finally, a task specific model head is applied.
@@
    GraphSAGE with Tabular ResNet is similar to the previous, with the tabular-level stage reducing the node embedding dimensionality, however, the operation is performed through a more sophisticated tabular ResNet model [gorishniy_revisiting_2021]. Notably, this model was previously used in [robinson2024relbench], allowing to directly align results between the RelBench and ReDeLEx benchmarks.
@@
    DBFormer is an implementation of the Transformer-based RDL model from [peleska_transformers_2024]. In constrast to the previous, [^6] The key difference can be viewed analogously to the “fusion” and “cooperation” in multi-modal learning [hu2021unit, liang2024foundations], considering the attributes as modalities. While the first two models fuse the representations of the discrete attributes at the beginning, the DBFormer allows for cooperation of the attributes through the GNN stage. the model retains the original node embedding dimensionality $h_{v}^{(l)}\in\mathbb{R}^{n\times d^{(l)}}$ while exploiting the attention mechanism [vaswani2017attention] for learning interactions between both the attributes and tuples through a custom message-passing scheme.
@@ 4.0.3 Classical Models
In addition to the selected RDL models, we include key representatives from related ML domains, including Gradient Boosted Decision Trees (GBDT; [natekin2013gradient]), Deep Tabular Learning (DTL; [borisov2022deep]) and Propositionalization (Prop.; [kramer2001propositionalization]). Particularly, we compare against the LightGBM [ke_lightgbm_2017]—representative of GBDT; the getML’s [getml] FastProp feature generator combined with XGBoost [chen2016xgboost]—representative of propositionalization; and the standalone tabular ResNet [gorishniy_revisiting_2021]—representative of deep tabular learning. Importantly, the LightGBM and the ResNet have access only to data from the task’s target table, as these models fall into the tabular learning category. In contrast, the propositionalization method of FastProp with XGBoost exploits the full RDB structure.
@@
Binary Classification (AUC ROC $\uparrow$)
rgastf1 orig. .576 .609 .917 .922 .535 .595 .900 .897 .895 .900 .904 .902
expenditures orig. .847 .852 .813 .813 .845 .847 .920 .917 .919 .915 .893 .892
geneea orig. .994 .989 .937 .953 .976 .981 .986 .983 .987 .975 .996 .994
temp. .990 .984 .945 .934 .980 .980 .986 .976 .990 .974 .998 .990
hepatitis orig. .628 .642 .964 .932 .698 .633 1.0 .997 1.0 .996 1.0 .999
imdb orig. .986 .986 .549 .546 .984 .985 .993 .993 .993 .993 .993 .993
mondial orig. .500 .500 NaN NaN .500 .528 .963 .954 .971 .920 .984 .941
movielens orig. .578 .613 .784 .806 .536 .600 .760 .789 .754 .791 .755 .788
musklarge orig. .500 .500 .619 .760 .452 .700 .876 .608 .900 .632 .905 .720
musksmall orig. .500 .500 .800 .778 .500 .583 .830 .856 .880 .878 .760 .756
mutagenesis orig. .901 .841 .917 .976 .845 .917 .933 .845 .940 .707 .955 .812
ncaa orig. .562 .475 .654 .791 .568 .543 .753 .736 .720 .746 .726 .712
studentloan orig. .500 .500 .728 .753 .500 .532 1.0 1.0 1.0 1.0 1.0 1.0
amazon ichurn .728 .730 NaN NaN .652 .653 .820 .824 .822 .825 .818 .821
uchurn .570 .571 NaN NaN .525 .531 .706 .707 .706 .707 .703 .704
avito clicks .561 .542 NaN NaN .536 .539 .651 .642 .653 .671 .647 .664
visits .536 .529 NaN NaN .514 .508 .698 .660 .699 .664 .694 .658
f1 dnf .675 .681 .666 .766 .624 .716 .703 .759 .726 .753 .702 .762
top3 .703 .745 .679 .851 .621 .748 .727 .830 .718 .831 .689 .782
stack badge .762 .740 .514 .513 .688 .674 .901 .890 .901 .891 .897 .885
engmt. .833 .823 .589 .590 .788 .781 .901 .905 .901 .907 .901 .904
trial s.out. .673 .692 .626 .699 .583 .665 .624 .676 .618 .692 .604 .656
Multiclass Classification (macro F1 score $\uparrow$)
accidents orig. .294 .292 .495 .492 .347 .357 1.0 1.0 1.0 1.0 1.0 1.0
temp. .166 .149 .319 .313 .193 .180 .816 .583 .819 .616 .820 .621
craftbeer orig. .406 .232 .006 .020 .394 .308 .369 .238 .483 .275 .476 .284
dallas orig. .520 .537 .550 .606 .460 .534 .440 .467 .502 .494 .394 .537
temp. .564 .348 .652 .347 .446 .236 .417 .411 .500 .377 .464 .360
diabetes orig. .190 .190 .402 .383 .189 .190 .866 .867 .888 .886 .880 .877
financial orig. .422 .416 .935 .886 .549 .471 .483 .451 .569 .621 .529 .578
temp. .495 .492 .466 .466 .214 .325 .292 .361 .376 .362 .378 .281
genes orig. .060 .060 .066 .100 .060 .060 1.0 .894 1.0 .941 1.0 .899
hockey orig. .626 .637 .689 .725 .653 .691 .692 .727 .697 .705 .697 .725
legalacts orig. .929 .921 .726 .723 .834 .823 .835 .823 .838 .827 .833 .823
temp. .852 .829 .727 .695 .757 .695 .787 .702 .778 .700 .775 .705
premiere orig. .270 .250 .630 .659 .417 .400 .594 .438 .524 .385 .452 .362
league temp. .250 .222 .626 .667 .467 .407 .531 .521 .572 .425 .467 .470
tpcd orig. .183 .176 .193 .196 .156 .168 .738 .743 .573 .579 .691 .693
webkp orig. .139 .130 .350 .366 .139 .130 .263 .238 .294 .209 .284 .273
Avg. Rank 4.49 3.61 4.71 2.75 2.66 2.79
2.91 2.41 3.16 1.53
@@ 4.1 Benchmarking tasks
We present comprehensive results over two types of node-level tasks—binary classification and multiclass classification. The tasks can be further differentiated by the origin of the target labels and usage of temporal values. Tasks performed on datasets from the RelBench collection use generated target table attributes, while tasks on datasets from the CTU Relational use existing target table attributes. Additionally, tasks from the CTU Relational can be both static and temporal, while tasks from the RelBench collection are always temporal. Static and temporal tasks differentiate based on the constrains forced upon the sampling algorithm while generating a sub-subgraph used for training the model, and by the method of splitting the dataset between the training, validation and test data. Static tasks use neighborhood sampling [hamilton2017inductive] constrained only by the maximum number of neighbors, and data splitting is carried out at random w.r.t. a given ratio (e.g. 70:15:15). In contrast, temporal tasks extend the neighborhood sampling by incorporating temporal constraints, assuming only directed edges from nodes with an older timestamp and, similarly, the splits are carried out w.r.t. the timestamps where all training entities must precede validation and testing data, respectively (see App. 0.A for the full experimental setup).
The results on classification tasks shown in Table 1 demonstrate a strong performance of the RDL models (Sec. 4.0.2) over the classical models (Sec. 4.0.3) on majority of the datasets. Specifically, the ResNet SAGE performs very well on binary classification tasks, especially on tasks with a training table generated from historical data. Nonetheless, both DBFormer and Linear SAGE perform just marginally worse compared to ResNet SAGE on average, showcasing the general robustness of the RDL paradigm. Notably, on the hepatits, mondial, student loan, accidents, and genes datasets, all the RDL models present near-perfect predictions while the classical models show an order of magnitude worse score, highlighting the contribution of the RDL representation.
@@
Base Joined Base Joined 2 layer Best
@@
ergastf1 orig. .576 .609 .913 .906 .535 .595 .873 .893 .900 .897 .904 .902
hepatitis orig. .628 .642 .920 .904 .698 .633 .696 .634 .998 .987 1.0 .997
mondial orig. .500 .500 .942 .855 .500 .528 .920 .917 .963 .935 .984 .941
movielens orig. .578 .613 .571 .613 .536 .600 .536 .600 .551 .618 .760 .789
ncaa orig. .562 .475 .742 .729 .568 .543 .664 .776 .672 .764 .753 .736
studentloan orig. .500 .500 1.0 1.0 .500 .532 1.0 1.0 1.0 1.0 1.0 1.0
amazon uchurn .570 .571 .561 .563 .525 .531 .525 .531 .699 .699 .706 .707
avito clicks .561 .542 .562 .550 .536 .539 .535 .538 .623 .647 .653 .671
visits .536 .529 .535 .532 .514 .508 .514 .509 .680 .655 .699 .664
stack badge .762 .740 .778 .754 .688 .674 .687 .675 .892 .880 .901 .890
@@
accidents orig. .294 .292 .369 .367 .347 .357 .359 .362 1.0 1.0 1.0 1.0
temp. .166 .149 .166 .152 .193 .180 .188 .177 .801 .671 .820 .621
craftbeer orig. .406 .232 .448 .271 .394 .308 .394 .308 .442 .335 .483 .275
diabetes orig. .190 .190 .190 .190 .189 .190 .189 .190 .871 .869 .888 .886
genes orig. .060 .060 .060 .060 .060 .060 .060 .060 1.0 .898 1.0 .894
premiere orig. .270 .250 .370 .306 .417 .400 .438 .349 .570 .409 .594 .438
league temp. .250 .222 .427 .339 .467 .407 .442 .402 .523 .419 .572 .425
tpcd orig. .183 .176 .822 .821 .156 .168 .671 .667 .600 .603 .738 .743
webkp orig. .139 .130 .139 .130 .139 .130 .139 .130 .233 .211 .294 .209
@@ 4.2 Tabular Learning
In this scenario, we compare results of the Tabular Learning (TL) models from the previous section to ones trained on new tables formed by join operations over the RDBs (Sec. 3.2.2). Particularly, we evaluate the TL models on tables generated by joins over both one-to-one and many-to-one relationships of the target table. [^7] This is similar to a recent RelGNN method [chen_relgnn_2025], albeit limited to the target table. Additionally, we include evaluation of RDL models with exactly $2$ layers in the graph neural stage, which is conceptually equivalent to the join operation. Finally, we include the overall best RDL models to put the results into context. Note that the previous Table 1 demonstrated that the TL methods perform significantly worse on an absolute majority of tasks. In this experiment, we aim to assess whether a simple RDB transformation could actually change the situation in some cases. Particularly, we select a subset of tasks from Table 1 where both the TL models showed at least $0.1$ worse score than the best RDL model. The results in Table 2 show that, indeed, on a number of datasets the TL models register a significant improvement with results sometimes comparable to the best of RDL. Notably, on the ergastf1, ncaa and tpcd they even set new best results. This experiment demonstrates existing weak spots in the new RDL approach [fey2024position], suggesting that caution and thorough analysis are still in order before deploying RDL on an RDB task.
@@ 4.3 Essential Characteristics
#Tab. 6 3 7 8 3 7 9 3.5 4 5
#FK 6 2 6 12 3 6 13 3.25 4 5
#Factual 35 18 26 48 26 47 110 23.25 28 32.75
#Cat. 11 3 11 18 14 17 88 10.25 13 14
#Num. 6 1 6 11 3 5 12 3 3 6.75
#Text 6 2 6 14 0 2 9 2.5 3 3
#Time 2 0 1 4 0 1 7 3.75 5 5
#Rows 472k 21k 1.5M 5.6M 11k 98k 546k 810k 1.4M 1.8M
#Links 752k 43k 2.2M 8.2M 32k 228k 1.1M 814k 1.7M 2.3M
Diameter 3 2 2 4 2 3 3 2.5 3 3
1-to-1 0 0 0 6 0 0 0 0 0 1.5
1-to-M 5 2 3 11 3 5 13 3.25 4 4.25
#Train 12k 0.8k 59k 654k 0.2k 0.5k 5k 0.5k 226k 455k
#T.Factual 5 1 5 13 2 5 6 4.25 12.5 20
T.Cat. 2 0 0 4 1 2 4 0.75 5.5 10
T.Num. 0 0 0 1 0 0 2 1.5 2.5 3
T.Text 1 0 1 4 0 0 3 1.5 2 2
T.Time 1 0 0 1 0 1 1 0.75 3 5
Eccentri. 10.64 7.92 12.24 17.84 9.36 13.90 19.20 4.59 7.74 9.92
Density 0.025 0.009 0.022 0.107 0.001 0.002 0.028 0.067 0.109 0.243
@@
The overarching aim of ReDeLEx is to assess common characteristics of RDBs and tasks in the context of a used learning approach. While a fully comprehensive assessment is out of scope of this short paper, in Table 3 we summarize characteristics of the databases against the respective performances of the various model types. Following the analysis, RDL models generally tend to perform well on datasets with a large number of training samples and links. Propositionalization achieves best results mostly on smaller datasets, yet with a large number of factual (non-key) columns. This is in line with some previous studies [Lavrač2021], despite the remaining prevalence of propositionalization methods in practice [getml]. The TL methods then tend to perform best when there is a higher number of factual columns in the target table, which aligns with natural intuition. Moreover, these allow capturing more diverse attribute types where, e.g., both LightGBM and deep TL models are capable of utilizing textual attributes.